OpenStudy (hawkfalcon):

Find the average rate of change for the following functions on the indicated intervals: f(x)=x3−2x;[0,4]

OpenStudy (anonymous):

? is this real

OpenStudy (hawkfalcon):

Yeah. Ikr:(

OpenStudy (hawkfalcon):

\[f(x) = x^3 - 2x; [0,4]\]

OpenStudy (anonymous):

average rate of change in (a,b) is: f(b)-f(a)/b-a

OpenStudy (anonymous):

rate of change = \[ \frac{f(4) - f(0)}{4 - 0} = \frac{4^3 - 2\times4 - \left( 0 - 0 \right)}{4} = \frac{64 - 8}{4} = 14 \]

OpenStudy (hawkfalcon):

O.o I like how they give us work which they haven't taught us. Thank you:)